The radius of a pipe 5cm and its thickness is 6mm. What is the external circumference of the pipe?
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

So, if 5cm is the INTERNAL radius, which I assume it is because they're giving you the thickness of the tube, then the external radius is 5.6cm or 56mm if you want to do it in mm
:
56*2*Pi = 112*Pi = external diameter of the tube in mm, divide the product by 10 if you want it in cm. Use your calculator.
John

Answer by ikleyn(52781)

.
This is an example on how a condition SHOULD NOT be formulated.